OpenStudy (anonymous):

What is a simpler form of the following expressions? (j + 7)(j – 7)

OpenStudy (abb0t):

FOIL

OpenStudy (agent0smith):

It's a difference of two squares: \[a^2 - b^2 = (a + b) (a-b)\]

OpenStudy (whpalmer4):

@ChanHamm Uh, no. \[(j+7)(j-7) = j(j-7) + 7(j-z) = j^2 -7j + 7j -49 = j^2 - 49\]